A non conducting spherical shell of  inner radius "a" and outer radius "b" has p= A/r where a is constant and r is the distance from the center of shell . in addition a positive point charge "q" is locatewd at center . What value should A have if electric field in the shell is to be uniform ?

Grade:12

1 Answers

Chetan Mandayam Nayakar
12 years ago

(∫(((A/r)4pi*r^2)/(4pikr^2))dr from r=a to r=b)=(∫((A/r))/(k)dr =Ak(ln(r/a))+(q/(4pikr^2)=constant,

so the derivative w.r.t. r=0